# A Simple yet Powerful Instance-Aware Prompting Framework for Training-free Camouflaged Object Segmentation

### v2

@@ Introduction
UCOS-DA (Zhang and Wu 2023) ICCV23 $U$ .755 .656 .085 .819 .689 .513 .086 .739 .700 .605 .127 .784
SdalsNet (Shou et al. 2025) AAAI25 $U$ .739 .642 .085 .824 .697 .525 .072 .780 .697 .601 .117 .799
EASE (Du et al. 2025) CVPR25 $U$ .800 .735 .056 .884 .773 .656 .040 .866 .749 .684 .098 .831
WS-SAM (He et al. 2023a) NeurIPS23 $P$ .813 .734 .057 .859 .790 .663 .039 .856 .718 .602 .102 .757
P-COD (Chen et al. 2024a) ECCV24 $P$ .822 .748 .051 .889 .784 .650 .042 .859 .798 .727 .074 .872
SAM-COD (Chen et al. 2024b) ECCV24 $P$ .858 .802 .041 .918 .831 .725 .031 .901 .820 .760 .066 .885
BoxSAM (Li et al. 2025a) IVC25 $P$ .810 - .069 .823 .808 - .042 .851 .745 - .102 .749
SS (Zhang et al. 2020) CVPR20 $S$ .723 .584 .090 .803 .684 .466 .068 .774 .684 .538 .119 .760
SCWS (Yu et al. 2021) AAAI21 $S$ .757 .659 .072 .846 .708 .535 .059 .813 .719 .616 .105 .810
CRNet (He et al. 2023b) AAAI23 $S$ .775 .688 .063 .855 .733 .576 .049 .832 .735 .641 .092 .815
WS-SAM (He et al. 2023a) NeurIPS23 $S$ .829 .757 .052 .886 .803 .680 .038 .877 .759 .667 .092 .818
MINet (Niu et al. 2024) ACM MM24 $S$ .793 .709 .061 .869 .749 .596 .049 .840 .750 .669 .091 .840
WSMD (Zha et al. 2024) AAAI24 $S$ .797 .700 .064 .868 .761 .600 .049 .839 .793 .704 .079 .866
SAM-COD (Chen et al. 2024b) ECCV24 $S$ .859 .795 .039 .912 .833 .728 .029 .904 .836 .779 .060 .903
BoxSAM (Li et al. 2025a) IVC25 $S$ .860 - .052 .894 .836 - .044 .883 .830 - .084 .860
SAM-COD (Chen et al. 2024b) ECCV24 $B$ .867 .813 .037 .923 .842 .745 .028 .914 .837 .786 .062 .901
BoxSAM (Li et al. 2025a) IVC25 $B$ .877 - .037 .925 .857 - .027 .919 .859 - .057 .908
LLaVA1.5+SAM (Liu et al. 2024) CVPR23 ZSZS .609 .543 .253 .676 .653 .551 .193 .719 .571 .526 .292 .631
CLIP_Surgey+SAM (Li et al. 2025b) PR25 ZSZS .712 .626 .152 .755 .661 .531 .174 .699 .671 .601 .179 .729
Grounded SAM (Ren et al. 2024) Arxiv24 ZSZS .771 .721 .135 .806 .745 .667 .146 .784 .683 .654 .203 .723
MMCPF (Tang et al. 2024) ACM MM24 ZSZS .768 .681 .082 - .733 .592 .065 - .749 .680 .100 -
GenSAM (Hu et al. 2024a) AAAI24 ZSZS .805 .745 .066 .863 .783 .682 .055 .845 .727 .648 .105 .799
ProMaC (Hu et al. 2024b) NeurIPS24 ZSZS .814 .762 .056 .884 .803 .710 .042 .875 .745 .685 .100 .830
RDVP-MSD (Yin et al. 2025) ACM MM25 ZSZS .823 .764 .056 .873 .825 .743 .038 .877 .796 .750 .081 .848
@@
Table 1: Quantitative comparison of COS methods across three benchmark datasets. The table compares state-of-the-art methods based on unsupervised, weakly-supervised, and training-free setups. ’$U$’ indicates unsupervised settings, while the weakly-supervised settings include ’$P$’ (point), ’$S$’ (scribble), and ’$B$’ (box). ’ZSZS’ represents zero-shot training-free approaches applied to COS. ’-’ denotes the absence of explicit mention and unavailable code in the corresponding paper. ’↑’ indicates higher is better, and ’↓’ indicates lower is better. The best results of different settings are highlighted in bold.
Existing training-free COS pipelines (Yin et al. 2025; Hu et al. 2024a, b; Tang et al. 2024) predominantly rely on a single semantic-level box prompt per image, which limits SAM (Kirillov et al. 2023) to return only one coarse mask, thereby hindering comprehensive segmentation when multiple camouflaged instances coexist. To overcome this bottleneck, the proposed Instance Mask Generator serves as the crucial bridge between the text prompts derived from the Text Prompt Generator and the segmentation masks output by the SAM. The process involves three frozen steps: (1) Box prompting: Grounding DINO (Liu et al. 2023) converts each foreground tag into multiple instance-level bounding box prompts. (2) Point prompting: Within every box prompt, the proposed Single-Foreground Multi-Background Prompting (SFMBP) strategy samples high-confidence foreground points and equally confident multi-background points, yielding instance-level point prompts. (3) Mask Prediction: Each instance-level visual prompt (i.e., a box–points pair) is fed to SAM for producing a mask for an instance; aggregating all masks yields a candidate instance mask that covers all camouflaged instances in the image.
Box Prompting
Existing training-free COS methods often fail to segment objects in scenarios containing multiple discrete camouflaged instances, as they typically provide a single bounding box per image for SAM to segment. According to SAM, accurate instance segmentation requires explicit object bounding box prompts (Kirillov et al. 2023). Therefore, the simplest solution to this problem is to provide an instance-level bounding box prompt for each camouflaged instance in the image. To this end, we introduce Grounding DINO (Liu et al. 2023) and leverage its instance-aware detection capabilities. Specifically, for each foreground tag $tag_{i}^{fg}$, we apply Grounding DINO to the pair $(X,tag_{i}^{fg})$ to generate multiple instance-level bounding box prompts $\mathcal{B}_{i}$:
$\mathcal{B}_{i}=\text{G-DINO}(X,tag_{i}^{fg}),$ (4)
where $\mathcal{B}_{i}$ represents the $N$ instance-level bounding box prompts generated by Grounding DINO for the $i$-th foreground tag $tag_{i}^{fg}$. This approach replaces the ”single box per image” assumption with detector-driven coverage of all camouflaged instances.
Point Prompting
RDVP-MSD (Yin et al. 2025) introduced the Region-Constrained (RC) strategy, where point prompts are sampled within a single box per image, thereby alleviating the issue that previous approaches (Hu et al. 2024a, b) mostly relied on image-level heatmaps to generate point prompts. However, this approach, like others, is still limited by two key factors: (1) the camouflage environment is often complex and diverse, making it difficult to generate sufficient background point prompts using a single background tag; (2) the bounding box is typically a single box per image, severely limiting its efficacy in scenarios containing multiple discrete camouflaged instances.
To mitigate these issues, we propose Single-Foreground Multi-Background Prompting (SFMBP) strategy, which leverages multi-background cues along with instance-level bounding box prompts to sample foreground and background points more effectively. Specifically, Spatial CLIP (Hu et al. 2024a) is used to generate a foreground heatmap $H^{f_{i}}$ and multiple background heatmaps $H^{b}=\{H^{b_{1}},H^{b_{2}},\dots,H^{b_{L}}\}$, where each $H^{b_{l}}$ corresponds to a specific background tag $tag^{b_{l}}$ from $\text{set}(tag^{bg})$, and $L$ is the length of $\text{set}(tag^{bg})$:
$H^{f_{i}}$ $=\text{Spatial CLIP}(X,tag_{i}^{fg}),$ (5)
$H^{b_{l}}$ $=\text{Spatial CLIP}(X,tag^{b_{l}}),\quad l=1,2,\dots,L,$ (6)
where $H^{f_{i}}$ is the foreground heatmap generated for the foreground tag $tag_{i}^{fg}$, and each $H^{b_{l}}$ is a background heatmap corresponding to a different background tag $tag^{b_{l}}\in\text{set}(tag^{bg})$.
These heatmaps encode pixel-level confidence scores for the foreground and background categories. By adaptively sampling high-confidence points from both the $i$-th foreground heatmap and all background heatmaps inside every instance-level bounding box $\mathcal{B}_{i,n}\in\mathcal{B}_{i}$, SFMBP ensures the acquisition of accurate instance-level point prompts, addressing both the issue of the complexity of diverse background environments and the single box limitation. Formally, the point sampling process can be described as follows:
$p_{i,n}^{fg}$ $=\left\{(x,y)\mid Norm(H^{f_{i}}\mid\mathcal{B}_{i,n})_{(x,y)}\geq 0.9\right\},$ (7)
$p_{i,n}^{\mathrm{bg}}$ $=\bigcup_{l=1}^{L}\left\{(x,y)\mid Norm(H^{b_{l}}\mid\mathcal{B}_{i,n})_{(x,y)}\geq 0.9\right\},$ (8)
where $p_{i,n}^{fg}$ represents the foreground point prompts inside the $n$-th box $B_{i,n}$ with high confidence, and $p_{i,n}^{\mathrm{bg}}$ represents the background point prompts inside the same box, sampled from the multiple background heatmaps. The $Norm(\cdot)$ denotes a normalization function, ensuring pixel-level confidence scores are normalized within a standard range(i.e., $[0,1]$).
Mask Prediction
[…]
$m_{i,n}=\text{SAM}(X\mid\langle\mathcal{B}_{i,n},\,p^{fg}_{i,n},\,p^{bg}_{i,n}\rangle),n=1,\dots,N,$ (9)
where the $N$ denotes the number of instance-level bounding boxes corresponding to the $tag_{i}^{fg}$. The resulting masks $\{m_{i,1},\ldots,m_{i,N}\}\in\{0,1\}^{H\times W}$ delineate all camouflaged instances associated with $tag^{fg}_{i}$. To yield a candidate instance mask that faithfully covers every instance, we stack them along a new dimension:
$mask_{i}=[m_{i,1};m_{i,2};...;m_{i,N}],$ (10)
where the $mask_{i}\in\{0,1\}^{N\times H\times W}$.
Self-consistency Instance Mask Voting
@@
Methods Venue Setting NC4K COD10K-TEST
@@
OSFormer (Pei et al. 2022) ECCV22 $F$ .444 .737 .451 .420 .713 .428
CLPNet (Cui, Sun, and Yuan 2023) ICIP23 $F$ .451 .748 .455 .433 .719 .432
QFormer (Dong et al. 2023) ACM MM23 $F$ .501 .768 .528 .454 .718 .479
DCNet (Luo et al. 2023) CVPR23 $F$ .528 .771 .565 .453 .707 .475
AQSFormer (Dong et al. 2024) ACM MM24 $F$ .505 .768 .535 .465 .738 .485
Mask2Camouflage (Phung and Shuai 2024) ACCV24 $F$ .538 .776 .583 .468 .725 .490
TPNet (He et al. 2024) ACM MM24 II .214 .483 .166 .183 .418 .143
@@
Following previous methods (Yin et al. 2025), we also adopt a self-consistency approach to select the most robust segmentation result. Specifically, we introduce a Self-consistency Instance Mask Voting (SIMV) mechanism to select the best instance segmentation mask rather than a semantic mask. In the SIMV mechanism, we need a semantic mask that is directly comparable across repeated runs. To achieve this, we convert the instance mask $mask_{i}$ into a semantic mask $M_{i}\in\{0,1\}^{H\times W}$ by performing a per-pixel logical OR $\bigvee$ over the instance dimension $N$:
$M_{i}(x,y)=\bigvee_{n=1}^{N}mask_{i}(n,x,y).$ (11)
We repeat the pipeline $I$ times (default $3$) with synonymic task-generic prompts $p_{i}^{g}$, resulting in a set of candidate semantic masks $\{M_{1},\dots,M_{I}\}$. To quantify their consistency, we apply the self-consistency rule:
$\hat{i}=\arg\min_{i}\;\left|M_{i}-\frac{1}{I}\sum_{i=1}^{I}M_{i}\right|.$ (12)
The output $M_{\hat{i}}$ and the corresponding $mask_{\hat{i}}$ represent the final COS and CIS predictions, respectively.
Experiments
Experiment Settings
Datasets and Evaluation Metrics
For COS, we evaluate on widely-used datasets, including NC4K ( (Lv et al. 2021), containing 4,121 images), COD10K-TEST ( (Fan et al. 2020), with 2,026 images), and CAMO-TEST ( (Le et al. 2019), consisting of 250 images). We use S-measure ($S_{\alpha}$(Fan et al. 2017), weighted F-measure ($F^{\omega}_{\beta}$), mean absolute error ($M$), and mean E-measure ($E_{m}^{\phi}$) to evaluate the performance. High-performing COS methods generally achieve higher values of $S_{\alpha}$, $F^{\omega}_{\beta}$, and $E_{m}^{\phi}$, along with a lower $M$ score. For CIS, we evaluate on NC4K (Lv et al. 2021) and COD10K-TEST (Fan et al. 2020) and report Average Precision ($AP$), $AP_{50}$, and $AP_{75}$. Higher values of $AP$, $AP_{50}$, and $AP_{50}$ indicate superior camouflaged instance segmentation performance.
Implementation details
In our experiments, we employ LLaVA-1.5-13B (Liu et al. 2024) as the MLLM. For the CLIP model, we use the CLIP (Radford et al. 2021) from the CS-ViT-L/14@336px variant. For the SAM, we deploy the HQ-SAM (Ke et al. 2023) built on the ViT-H version. Our method is entirely train-free, requiring no fine-tuning. All experiments are conducted using the PyTorch framework, and we utilize two NVIDIA GeForce RTX 3090 GPUs with 24 GB of memory each for evaluation. The reported results are based on a single execution without repeated trials or statistical variation analysis.
Comparison with State-of-the-Art Methods
Quantitative and Qualitative Comparison
We compare the proposed IAPF with unsupervised, weakly-supervised, and training-free methods on the NC4K, COD10K-TEST, and CAMO-TEST datasets, as shown in Table 1. In the zero-shot training-free setting, our method achieves significant improvements over all compared methods, including higher $S_{\alpha}$, $F_{\beta}^{\omega}$, and $E_{m}^{\phi}$ values. This performance improvement is mainly due to the proposed stance Mask Generator, which generates multiple instance-level bounding boxes, and the SFMBP strategy. These innovations enable more accurate segmentation, especially in challenging scenarios with multiple camouflaged instances. Compared to unsupervised and weakly-supervised methods, the proposed IAPF achieves competitive performance without any form of supervision or training.
Qualitative results in Figure 3 show that, as the number of camouflaged instances increases, previous training-free COS methods struggle to segment the objects accurately, while our method maintains strong performance across all scenarios.
Evaluation on Camouflaged Instance Segmentation
We further evaluate our method on the Camouflaged Instance Segmentation (CIS) task, reporting quantitative results on the NC4K and COD10K-TEST datasets in Table 2. Compared to recent fully-supervised (’$F$’) and image-level supervised (’$I$’) CIS approaches, our method achieves competitive or even superior performance, notably attaining the highest $AP_{50}$ scores on both NC4K and COD10K-TEST datasets. Remarkably, our results are obtained entirely in a training-free manner using only a single task-generic prompt, underscoring the effectiveness of the instance-aware prompting strategy for accurate instance segmentation without reliance on costly annotations.

### v3

@@ 3.3.2 Point Prompting
UCOS-DA [44] ICCV23 $U$ .755 .656 .085 .819 .689 .513 .086 .739 .700 .605 .127 .784
SdalsNet [34] AAAI25 $U$ .739 .642 .085 .824 .697 .525 .072 .780 .697 .601 .117 .799
EASE [7] CVPR25 $U$ .800 .735 .056 .884 .773 .656 .040 .866 .749 .684 .098 .831
WS-SAM [13] NeurIPS23 $P$ .813 .734 .057 .859 .790 .663 .039 .856 .718 .602 .102 .757
P-COD [1] ECCV24 $P$ .822 .748 .051 .889 .784 .650 .042 .859 .798 .727 .074 .872
SAM-COD [2] ECCV24 $P$ .858 .802 .041 .918 .831 .725 .031 .901 .820 .760 .066 .885
×BoxSAM [23] IVC25 $P$ .810 - .069 .823 .808 - .042 .851 .745 - .102 .749
SS [43] CVPR20 $S$ .723 .584 .090 .803 .684 .466 .068 .774 .684 .538 .119 .760
SCWS [41] AAAI21 $S$ .757 .659 .072 .846 .708 .535 .059 .813 .719 .616 .105 .810
CRNet [14] AAAI23 $S$ .775 .688 .063 .855 .733 .576 .049 .832 .735 .641 .092 .815
WS-SAM [13] NeurIPS23 $S$ .829 .757 .052 .886 .803 .680 .038 .877 .759 .667 .092 .818
MINet [29] ACM MM24 $S$ .793 .709 .061 .869 .749 .596 .049 .840 .750 .669 .091 .840
WSMD [42] AAAI24 $S$ .797 .700 .064 .868 .761 .600 .049 .839 .793 .704 .079 .866
SAM-COD [2] ECCV24 $S$ .859 .795 .039 .912 .833 .728 .029 .904 .836 .779 .060 .903
×BoxSAM [23] IVC25 $S$ .860 - .052 .894 .836 - .044 .883 .830 - .084 .860
SAM-COD [2] ECCV24 $B$ .867 .813 .037 .923 .842 .745 .028 .914 .837 .786 .062 .901
×BoxSAM [23] IVC25 $B$ .877 - .037 .925 .857 - .027 .919 .859 - .057 .908
LLaVA1.5+SAM [25] CVPR23 ZSZS .609 .543 .253 .676 .653 .551 .193 .719 .571 .526 .292 .631
CLIP_Surgey+SAM [24] PR25 ZSZS .712 .626 .152 .755 .661 .531 .174 .699 .671 .601 .179 .729
Grounded SAM [33] Arxiv24 ZSZS .771 .721 .135 .806 .745 .667 .146 .784 .683 .654 .203 .723
×MMCPF [35] ACM MM24 ZSZS .768 .681 .082 - .733 .592 .065 - .749 .680 .100 -
GenSAM [16] AAAI24 ZSZS .805 .745 .066 .863 .783 .682 .055 .845 .727 .648 .105 .799
ProMaC [17] NeurIPS24 ZSZS .814 .762 .056 .884 .803 .710 .042 .875 .745 .685 .100 .830
×UpGen [8] TIP25 ZSZS .838 .788 .054 .887 .806 .708 .051 .856 .779 .718 .091 .833
×INT [18] IJCAI25 ZS - - - - .808 .722 .037 .883 .772 .734 .086 .853
RDVP-MSD [40] ACM MM25 ZSZS .823 .764 .056 .873 .825 .743 .038 .877 .796 .750 .081 .848
@@
Table 1: Quantitative comparison of COS methods across three benchmark datasets. The table compares state-of-the-art methods based on unsupervised, weakly-supervised, and training-free setups. ’$U$’ indicates unsupervised settings, while the weakly-supervised settings include ’$P$’ (point), ’$S$’ (scribble), and ’$B$’ (box). ’ZSZS’ represents zero-shot training-free approaches applied to COS. ’×’ denotes the unavailable code in the corresponding paper. ’↑’ indicates higher is better, and ’↓’ indicates lower is better. The best results of different settings are highlighted in bold.
3.3.3 Mask Prediction
Given the $i$-th repetition’s foreground tag $tag^{fg}_{i}$ and its enumerated boxes, SFMBP yields an instance-level prompt triplet for each detected instance, $\langle\mathcal{B}_{i,n},\,p^{fg}_{i,n},\,p^{bg}_{i,n}\rangle_{n=1}^{N}$—one box with its foreground and multi-background point prompts. Unlike prior semantic-level prompting that feeds SAM a single global/region cue, these per-instance triplets preserve instance identity and prevent mask bleeding across neighboring camouflaged objects. Each triplet is forwarded to the frozen SAM to obtain a camouflaged instance:
$m_{i,n}=\text{SAM}(X\mid\langle\mathcal{B}_{i,n},\,p^{fg}_{i,n},\,p^{bg}_{i,n}\rangle),\,n=1,\dots,N$ (10)
where $N$ is the number of instance-level box prompts associated with $tag^{fg}_{i}$. The resulting set $\{m_{i,1},\ldots,m_{i,N}\}\subset\{0,1\}^{H\times W}$ delineates the $N$ camouflaged instances discovered at repetition $i$.
To retain instance granularity for downstream use (e.g., per-instance analysis and subsequent self-consistency voting), we stack the masks along a new dimension to form a candidate instance mask:
$mask_{i}=[m_{i,1};m_{i,2};\ldots;m_{i,N}],$ (11)
with $mask_{i}\in\{0,1\}^{N\times H\times W}$.
3.4 Self-consistency Instance Mask Voting
In the fully frozen setting, using synonymous prompts $p_{i}^{g}$ ($i=1,\dots,I$) leads to MLLM-induced variability: the frozen MLLM may return slightly different captions/tags across synonym choices and decoding passes (prompt sensitivity and stochastic decoding), which in turn perturbs the detector’s box prompts and the SFMBP’s point prompts. For fair comparability across repetitions, we first project each candidate instance mask $mask_{i}\in\{0,1\}^{N\times H\times W}$ to a semantic mask $M_{i}\in\{0,1\}^{H\times W}$ by performing a per-pixel logical OR $\bigvee$ over the instance dimension $N$:
$M_{i}(x,y)\;=\;\bigvee_{n=1}^{N}mask_{i}(n,x,y).$ (12)
Running the pipeline $I$ times (default $3$) produces $\{M_{1},\dots,M_{I}\}$. We then select the repetition that is most self-consistent with the cohort:
$\hat{i}\;=\;\arg\min_{i}\;\left\|\,M_{i}\;-\;\frac{1}{I}\sum_{j=1}^{I}M_{j}\,\right\|_{2},$ (13)
where $\|\cdot\|_{2}$ denotes the pixel-wise $\ell_{2}$ distance over binary masks. Finally, $M_{\hat{i}}$ and the corresponding $mask_{\hat{i}}$ are reported as the COS and CIS outputs, respectively—voting is conducted on the semantic projection for comparability, while the selected index $\hat{i}$ returns the full instance-aware solution.
4 Experiments
4.1 Experiment Settings
4.1.1 Datasets and Evaluation Metrics
We evaluate COS on NC4K [28] (4,121), COD10K-TEST [10] (2,026), and CAMO-TEST [22] (250) using Structure-measure $S_{\alpha}$ [9], weighted F-measure $F^{\omega}_{\beta}$, mean E-measure $E^{\phi}_{m}$, and MAE $M$ (higher is better except $M$). For CIS, we follow COCO-style metrics and report $AP$, $AP_{50}$, and $AP_{75}$ on NC4K and COD10K-TEST. For downstream applications, we additionally evaluate—still in the training-free setting—on ACOD-12K-TEST (1,492; concealed crop detection; RGB-D, using RGB only) [37] and PlantCamo (1,250; plant camouflage) [38], evaluated with the same COS metrics as above.
4.1.2 Implementation details
In our experiments, we employ LLaVA-1.5-13B [25] as the MLLM. For the detector-agnostic box enumerator, we use the Grounding DINO [26]. For the CLIP model, we use the CLIP [32] from the CS-ViT-L/14@336px variant. For the SAM, we deploy the HQ-SAM [20] built on the ViT-H version. Our method is entirely train-free, requiring no fine-tuning. All experiments are conducted using the PyTorch framework, and we utilize two NVIDIA GeForce RTX 3090 GPUs with 24 GB of memory each for evaluation.
4.2 Comparison with State-of-the-Art Methods
4.2.1 Quantitative and Qualitative Comparison
We compare the proposed IAPF with unsupervised, weakly-supervised, and training-free methods on the NC4K, COD10K-TEST, and CAMO-TEST datasets, as shown in Table 1. In the zero-shot training-free setting, our method achieves significant improvements over all compared methods, including higher $S_{\alpha}$, $F_{\beta}^{\omega}$, and $E_{m}^{\phi}$ values. IAPF exceeds unsupervised baselines and, on NC4K and COD10K, even surpasses weakly supervised methods, all without any supervision or finetuning.
Qualitative results in Figure 3 show that, as the number of camouflaged instances increases, previous training-free COS methods [16, 17, 40] struggle to segment the objects accurately, while our method maintains strong performance across all scenarios.
4.2.2 Evaluation on CIS datasets
We further evaluate our method on the Camouflaged Instance Segmentation (CIS) task, reporting quantitative results on the NC4K and COD10K-TEST datasets in Table 2. Compared to recent fully-supervised (’$F$’) and image-level supervised (’$I$’) CIS approaches, our method achieves competitive or even superior performance, notably attaining the highest $AP_{50}$ scores on both NC4K and COD10K-TEST datasets. Remarkably, our results are obtained entirely in a training-free manner using only a single task-generic prompt, underscoring the effectiveness of the instance-aware prompting strategy for accurate camouflaged instance segmentation without reliance on costly annotations.
@@ 4.2.2 Evaluation on CIS datasets
4.3 Ablation Study
4.3.1 Effectiveness of the Modules
We perform ablation studies to examine the effectiveness of each key component in the IAPF framework: Text Prompt Generator (TPG), Grounding DINO (G), Single-Foreground Multi-Background Prompting (SFMBP), and Self-consistency Instance Mask Voting (SIMV). As shown in Table 3, removing TPG (setting $1$), G and SFMBP (settings 2.12.1-$2.3$), or SIMV (setting $3$) clearly degrades performance, verifying the necessity of each module. Moreover, comparisons between settings ($2.1$ vs. $2.3$) and ($2.2$ vs. $4$) further demonstrate that introducing Grounding DINO not only grants the IAPF the critical instance-awareness capability but also significantly improves overall segmentation accuracy by allowing the model to localize more camouflaged instances. The complete method (setting $4$) consistently outperforms all variants on both NC4K and COD10K-TEST datasets, confirming that each proposed component contributes distinctly and effectively to the final COS performance.
4.3.2 Effectiveness of Repetition Number
Methods Setting NC4K COD10K-TEST
@@ 4.3.2 Effectiveness of Repetition Number
OSFormer [30] $F$ .444 .737 .451 .420 .713 .428
CLPNet [4] $F$ .451 .748 .455 .433 .719 .432
QFormer [5] $F$ .501 .768 .528 .454 .718 .479
DCNet [27] $F$ .528 .771 .565 .453 .707 .475
AQSFormer [6] $F$ .505 .768 .535 .465 .738 .485
Mask2Camouflage [31] $F$ .538 .776 .583 .468 .725 .490
TPNet [15] II .214 .483 .166 .183 .418 .143
@@
4.3.3 Performance Comparison of Different Camouflaged Instance Counts
The proposed IAPF maintains consistently high accuracy as instance density grows. The larger percentage swing in $M$ for all methods is expected because the absolute MAE values are very small; nonetheless, the absolute MAE of the IAPF remains the lowest. These results confirm that the instance-aware prompting pipeline––particularly the detector-agnostic box enumerator–based box prompts and SFMBP point sampling––provides strong resilience to cluttered scenes with many camouflaged objects.
4.3.4 Comparison With Other SOTAs on a Fair Basis
To ensure a fair comparison, we evaluate all methods using the same model configuration, including LLaVA-1.5-13B as the MLLM, CS-ViT-L/14@336px for CLIP, and HQ-SAM built on ViT-H, as mentioned in Section 4.1.2. The results in Table 6 demonstrate that our method outperforms GenSAM [16] and ProMaC [17] by a significant margin. RDVP-MSD [40], which maintains the same configuration as our approach, shows no performance improvement, confirming the stability of our setup. Even with identical model configurations, our method consistently leads in all metrics, clearly showcasing its superior ability to handle multi-instance camouflage and produce accurate, training-free segmentation.
4.4 Downstream Applications
We evaluate on two real-world, multi-instance datasets—ACOD-12K (dense agricultural crop, RGB-only) and PlantCamo (plant camouflage). In these scenes, targets are small and dispersed; other training-free baselines that feed SAM semantic-level prompts collapse multiple objects or leak to textured backgrounds, while IAPF preserves instance separation (Figure 4). Quantitatively, IAPF achieves the best results across all metrics on both datasets, with clear margins (see Table 7), demonstrating that upgrading prompts from semantic-level to instance-level is crucial for robust generalization to downstream tasks with multiple scattered objects.